# Underactuated Control of Multiple Soft Pneumatic Actuators via Stable Inversion

### v1

@@ III-A Mechatronic Design
The mechatronic system design can be illustrated in Fig. 2 (a), (b), and (c). The soft gripper is composed of two main components such as soft fingers and a syringe pump as in Fig. 1. The design methodology of each component will be elaborated in the following paragraphs.
@@ III-A1 Soft Actuator Design
The soft actuator is designed under an optimal model-based design framework which considers force/torque, bendability, and controllability simultaneously during the design stage [21]. The dimensional parameters of a soft pneumatic actuator is displayed in Fig. 2 (a), which is the cross-sectional view of the soft actuator’s chamber room. The optimal dimensional parameters are searched by the optimization framework below
@@
where $a$ is the top of the chamber to the neutral surface, $b$ is the neutral surface to the bottom of the chamber, $a+b$ is the height of the soft actuator as in Fig. 2 (a), $E$ is Young’s modulus of the selected material, and $n$ is a parameter related to soft materials and determined by experiments [25]. Note that $w$ and $t$ represent the width and wall thickness of the cross-sectional area (Fig. 2 (a)). However, they usually hit the upper and lower bounds respectively, so they are not included in the (LABEL:eqn_opt) and determined by the designer.
The $\bar{T}(p)$ represents the Pressure-to-Force/Torque model which is obtained by mechanics analysis of the soft actuator [21], while the $\bar{\theta}(p)$ stands for the Pressure-to-Bending model which is derived by a nonlinear mechanics theory [25, 21]. The constraint of $Ew(a+b)^{n}$ aims to place the natural frequency of the soft actuator in the desired range. The remaining parameters which are not considered in the (LABEL:eqn_opt) include the Young modulus, length of the structure, and number of chamber rooms. They will be discussed in the following paragraph.
The range of dimensional parameters is selected by referencing the size of human fingers [26, 27], so the constraint of $a$, $b$, and the value of $w$ are determined and $n$ is decided by the selected soft material. To position the natural frequency in the desired range (2 - 3 $rad/s$), the Smooth-on Ecoflex®Dragon Skin 20 is selected and its Young’s modulus is 0.34 $MPa$. The length of the soft actuator and the number of chambers are couple. The more the number of chambers, the longer the length. The length of 100 $mm$ is chosen to avoid the buckling effect caused by the long structure and the corresponding number of the chamber rooms is 6.
The soft actuator is fabricated by two molds as illustrated in Fig. 2 (b). There are upper and bottom components on the left side of Fig. 2 (b). The Ecoflex®Dragon Skin 20 is in the liquid state, and its curing time is around 4 hours. A flex sensor is embedded into the bottom component as shown Fig. 2 (b) before the liquid rubber becomes a solid state. When the two components are removed from the molds, they are bonded by the silicone adhesive Smooth-on Sil-poxy®, as top right of Fig. 2 (b). The appearance of the soft actuator is demonstrated in bottom right of Fig. 2 (b).
@@ III-A2 Syringe Pump Design
The precision of the linear actuator and the volume of the syringe have an influence on the accuracy and controllability of the syringe pump. The linear actuator, Fulride by NSK Ltd., and a syringe with 150 $mL$ are chosen to fabricate the syringe pump. The accuracy of the Fulride could be $\mu m$ scale and the volume of the syringe could provide pressurize up to three soft actuators to generate $\pi/2$ rads. Some custom-made components are manufactured by 3D printers in order to assemble the syringe and the linear actuator.
@@ III-A3 Multi-finger Soft Gripper
Multiple soft pneumatic actuators and the syringe pump form a soft gripper module, which is assembled by 3D printed connectors and rubber tubes. The detailed compositions of the soft gripper including sensor setup will be described in Sec. IV-C.
@@ III-B System Modeling
Prior to designing the algebraic controller for the soft gripper, we need the full system dynamical model of both the soft actuators and the syringe pump. The models of soft actuators and the syringe pump are cascaded to obtain the full system model matrix. The model of each component is built in the following subsection.
@@ III-B1 Modeling Soft Actuators
The dynamical model of the soft pneumatic actuator is obtained by modeling its approximated structure as Fig. 2 (d). Due to the large bending nature of the soft actuator, a nonlinear second-order model is utilized to model and capture its motions. The nonlinear second-order model is described as [28]
@@
where $\theta$ is the bending as Fig. 2 (e), $F(p)$ is the equivalent force generated by the input pressure $p$ from the syringe pump, $M_{eq}$ is the equivalent mass of the soft actuator, $C_{n}$ is the damping constant of the soft actuator. The $M_{eq}$ and $C_{n}$ currently are estimated by system identification of the true responses in MATLAB, which has an average fitting accuracy of approximately $95.3\%$. The $K_{n}$ is the spring constant displayed as
@@
where $L_{0}$ is the initial length of the soft actuator, $n$ is the same as the parameter in (LABEL:eqn_opt) and it determined by either experiments or data-driven approach [28], $I_{n}$ is the modified moment of inertia for a large deflection component [25] and it is displayed as
@@
If $n=1$, the dynamical equations (7), (9), and (8) become linear equations. Given that the linear model is valid from $\theta=0$ to $94\pi/$ 9 rads [28], we use $n=1$ in this research. The (LABEL:eqn_opt) is rewritten as
@@
The state-space form is therefore written as
@@
where $F(P)$ is represented as $c\cdot p$, $c$ is a constant affected by $a$, $b$, $w$, and $t$ [23]. If we stack the system model of $n$ fingers, the state-space form becomes
$A_{stk}=\begin{bmatrix}A_{1}&0&0\\ 0&\ddots&0\\ 0&0&A_{n}\end{bmatrix},B_{stk}=\begin{bmatrix}B_{1}\\ \vdots\\ B_{n}\end{bmatrix},C_{stk}=\begin{bmatrix}C_{1}\\ \vdots\\ C_{n}\end{bmatrix}^{T}$ (12)
According to (4), we get the system equation in the Laplace domain. Thus, the system model matrix $P(s)^{n\times 1}$ is described as
${P}(s)=\begin{bmatrix}\frac{c\cdot p/M_{eq\_1}}{s^{2}+(C_{n\_1}/M_{eq\_1})s+K_{n\_1}/M_{eq\_1}}\\ \vdots\\ \frac{c\cdot p/M_{eq\_n}}{s^{2}+(C_{n\_n}/M_{eq\_n})s+K_{n\_n}/M_{eq\_n}}\end{bmatrix}$ (13)
Although the two fingers have the same dimensional parameters such as height, weight, etc., their $C_{n}$ and $K_{n}$ in (10) are slightly different due to fabrication errors and uncertainty of soft materials.
@@
This nature leads to asynchronized motions when a feedback controller is applied. The asynchronized motions further lead to grasping failure[7]. This study addressed this issue by algebraic control with a single syringe pump and the experimental results will be disclosed in Sec. IV-E.
@@ III-B2 Modeling Syringe Pump
The configuration of the syringe pump can be visualized in Fig. 2 (f). The dynamical modeling of the syringe pump starts at the linear motor. When the linear motor works, it will move the slider
@@
$\begin{split}Q_{i}=Av_{s}=\frac{l}{2\pi}\omega_{m}\end{split}$ (16)
@@
where $A$ is the inner cross-sectional area of the syringe, and $Q_{i}$ is the output air flow rate of the syringe. When $Q_{i}$ is divided by the capacity of the soft actuator, it becomes the pressure-changing rate inside the soft actuator
$\begin{split}\dot{p}=\frac{Q_{i}}{C_{i}}=\frac{Al}{2\pi C_{i}}\omega_{m}\end{split}$ (17)
@@
The dynamics of the syringe pump is the first-order system. The $\omega_{m}$ has a speed limit of 5 $rev/s$. The $C_{i}$ will expand as it is pressurized; however, its effect can be ignored as the input pressure is below 0.1 $MPa$ and the bending angle is below $32\pi/$ 3 $rad$. The $C_{i}$ here is considered as a constant.
The full model of a single soft actuator is the cascade of the (10) and (17).
$A_{1}=\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&-\frac{K_{n}}{M_{eq}}&-\frac{C_{n}}{M_{eq}}\end{bmatrix},B_{1}=\begin{bmatrix}0\\ 0\\ 1\end{bmatrix},C_{1}=\begin{bmatrix}\frac{c\cdot pAl}{2\pi C_{i}M_{eq}}\\ 0\\ 0\end{bmatrix}^{T}$ (18)
@@
The system is third-order and has a pole at the imaginary axis that makes the system marginally stable. Hence, the full system model matrix $P(s)^{n\times 1}$ in Laplace domain can be obtained by referencing (4) and (12)
${P(s)}=\begin{bmatrix}\frac{c\cdot pAl/2\pi C_{i}M_{eq\_1}}{s^{3}+(C_{n\_1}/M_{eq\_1})s^{2}+(K_{n\_1}/M_{eq\_1})s}\\ \vdots\\ \frac{c\cdot pAl/2\pi C_{i}M_{eq\_n}}{s^{3}+(C_{n\_n}/M_{eq\_n})s^{2}+(K_{n\_n}/M_{eq\_n})s}\end{bmatrix}$ (19)
@@
The full system matrix is causal and the minimum phase and the $P(s)$ is full rank and invertible.
@@ III-C Controller Design and Analysis
The full system matrix $P(s)$ will be utilized to design the algebraic controller. The system equation can be written as the (5) given the desired $Y(s)_{d}^{n_{y}\times 1}$ vector. The system equation is rewritten as
$\begin{bmatrix}Y_{d1}(s)\\ \vdots\\ Y_{dn}(s)\end{bmatrix}=\begin{bmatrix}P_{1}(s)\\ \vdots\\ P_{n}(s)\end{bmatrix}{U(s)}$ (20)
where $Y_{di}(s)$ represents the desired output of $P_{i}(s)$, and $1=1,...,n$. The stable inversion is composed of the feedforward and feedback loop [14]. The feedforward controller is obtained by solving (5) to get $U(s)$ and the additional feedback loop aims to address the system perturbation of the mechatronic system as shown in Fig. 3.
Some theorems are introduced and will be implemented to design the algebraic control including the feedforward control and feedback loop.
@@
(section III-B,  [14]) Let $P(s)$ be non-square ($rank_{\mathbb{R}(s)}(P)=n_{u}<n_{y}$), then there exists an $P\textsuperscript{\textdagger}(s):=(P^{*}(s)P(s))^{-1}P^{*}(s)$ satisfying $P\textsuperscript{\textdagger}(s)P(s)=I$. Thus, an approximate solution $U^{a}(s)$ is defined as
@@
4) $U^{a}(s)\in\mathcal{RH}_{\infty}$ ($\mathcal{RH}_{\infty}$ represents the Hardy space)
@@
(section III-B, [14]) Let $\omega_{cl}$ denote the bandwidth (BW) of the system. So the filter is defined as
$H_{i}(j\omega)\approx I\Leftrightarrow~{}\omega\ll\omega_{cl}$
$H_{i}(j\omega)\approx 0\Leftrightarrow~{}\omega\gg\omega_{cl}$
$H_{i}(j\omega)\not\approx\{0,I\}\Leftrightarrow~{}\omega$ close to $\omega_{cl}$
Since the bandwidth of the system is $\omega_{cl}$, the filter is designed by the Definition 1. The $H(s)$ of this mechatronic system is designed as a first-order low-pass filter taking the form $\frac{b}{s+a}$, considering its simplicity and stability.
In practice, the system has perturbation or the true system varies with the mathematical system. The system perturbation leads to unexpected dynamics which degrades the performance of the designed controller. Next, we consider the perturbed plant to be a member of all possible plants
$\Pi\in\{(I+{\Delta}{W_{T}})P~{}|~{}\forall\parallel\Delta\parallel_{\infty}\leq 1\}$ (22)
where the transfer function matrix $W_{T}\in\mathcal{RH}_{\infty}$ displays the spatial and frequency characteristics of the uncertainty. The symbol $\Delta$ denotes an unknown yet stable function, bounded within a specific norm [29]. A general approach to defining the robustness weight function $W_{T}$ is described below [30].
$\left|\frac{M_{ik}e^{j\phi_{ik}}}{M_{i}e^{j\phi_{i}}}-1\right|\leq\left|W_{T}(j\omega_{i})\right|,i=1,\ldots,m;k=1,\ldots,o$ (23)
The magnitude and phase values are assessed over a range of frequencies, denoted as $\omega_{i}$ (ranging from $i=1$ to $m$), and the experiment is repeated $n=5$ times. The notation ($M_{ik},\phi_{ik}$) refers to the magnitude-phase measurements corresponding to frequency $\omega_{i}$ and the $k$th experiment iteration. Similarly, ($M_{i},\phi_{i}$) represents the magnitude-phase pairs for the nominal plant $T$.
The error system (with perturbed term) can be compensated by using the output feedback as displayed in Fig. 3. The $u_{ff}$ is calculated based on the Theorem 2. It is assumed that the output of the real (uncertain) system can be measured such as sensors or observers. Since the nominal system output can be computed with the combined input $u_{c}$, we have the output difference $y_{\Delta}(t)$. With this output difference, we can compensate for the error by the following theorem.

### v2

@@ III-A Mechatronic Design
The mechatronic system design is illustrated in Fig. 2 (a), (b), and (c). The soft gripper is composed of two main components: soft fingers and a syringe pump, as in Fig. 1. The design methodology of each component will be elaborated in the following paragraphs.
@@ III-A1 Soft Actuator Design
The soft actuator is designed under an optimal model-based design framework that considers force/torque, bendability, and controllability simultaneously during the design stage [26]. The dimensional parameters of a soft pneumatic actuator are as shown in Fig. 2 (a), the cross-sectional view of the soft actuator’s chamber room. The optimal dimensional parameters are searched by the optimization framework below
@@
where $a$ is the top of the chamber to the neutral surface, $b$ is the neutral surface to the bottom of the chamber, $a+b$ is the height of the soft actuator as in Fig. 2 (a), $E$ is Young’s modulus of the selected material, and $n$ is a data-driven parameter related to soft materials [29]. Note that $w$ and $t$ represent the width and wall thickness of the cross-sectional area (Fig. 2 (a)). However, the optimal values for these two parameters tend to reach their respective upper and lower bounds, so they are not included in the (LABEL:eqn_opt) and are determined by the designer.
The $\bar{T}(p)$ represents the Pressure-to-Force/Torque model, which is obtained by mechanics analysis of the soft actuator [26], while the $\bar{\theta}(p)$ stands for the Pressure-to-Bending model which is derived by a nonlinear mechanics theory [29, 26]. Both $\bar{T}(p)$ and $\bar{\theta}(p)$ are functions of the dimensional parameters, $a$, $b$, $w$, and $t$ as shown in Fig. 2 (a). There exists an optimal parameter set that maximizes the objective function of (LABEL:eqn_opt), namely $\bar{T}(p)+\bar{\theta}(p)$ [26]. The parameter set is searched by optimization algorithms. The constraint of $Ew(a+b)^{n}$ aims to place the natural frequency of the soft actuator in the desired range. The remaining parameters that are not considered in the (LABEL:eqn_opt) include the Young modulus, length of the structure, and number of chamber rooms. They will be discussed in the following paragraph.
The range of dimensional parameters is selected by referencing the size of human fingers [30], so the constraint of $a$, $b$, and the value of $w$ are determined and $n$ is decided by the selected soft material. To position the natural frequency in the desired range (2 - 3 $rad/s$), the Smooth-on Ecoflex® Dragon Skin 20 is selected and its Young’s modulus is 0.34 $MPa$. The length of the soft actuator and the number of chambers are coupled. The more the number of chambers, the longer the length. The length of 100 $mm$ is chosen to avoid the buckling effect caused by the long structure and the corresponding number of the chamber rooms is 6.
The soft actuator is fabricated by two molds as illustrated in Fig. 2 (b). There are upper and bottom components on the left side of Fig. 2 (b). The Smooth-on Ecoflex® Dragon Skin 20 is in the liquid state, and its curing time is around 4 hours. A flex sensor is embedded into the bottom component, as shown in Fig. 2 (b), before the liquid rubber becomes a solid state. When the two components are removed from the molds, they are bonded by the silicone adhesive Smooth-on Sil-poxy®, as shown in the top right of Fig. 2 (b). The appearance of the soft actuator is as shown in the bottom right of Fig. 2 (b).
@@ III-A2 Syringe Pump Design
The precision of the linear actuator and the volume of the syringe have an influence on the accuracy and controllability of the syringe pump. The linear actuator, Fulride by NSK Ltd., and a syringe with 150 $mL$ are chosen to fabricate the syringe pump. The accuracy of the Fulride could be $\mu m$ scale and the volume of the syringe could provide pressurize up to three soft actuators to generate $\pi/2$ $rads$. Some 3D-printed components, which are easy to make and lightweight, are designed and manufactured to assemble the syringe and the linear actuator.
@@ III-A3 Multi-finger Soft Gripper
Multiple soft pneumatic actuators and the syringe pump form a soft gripper module, which is assembled by 3D-printed connectors and rubber tubes. The detailed compositions of the soft gripper, including sensor setup, will be described in Sec. IV-C.
@@ III-B System Modeling
Prior to designing the underactuated controller for the soft gripper, we need the full system dynamical model of both the soft actuators and the syringe pump. The models of soft actuators and the syringe pump are cascaded to obtain the full system model matrix. The model of each component is developed in the following subsection.
@@ III-B1 Modeling Soft Actuators
The dynamical model of the soft pneumatic actuator is obtained by modeling its approximated structure as shown in Fig. 2 (d). Due to the large bending nature of the soft actuator, a nonlinear second-order model is utilized to model and capture its motions. The nonlinear second-order model is described as [31]
@@
where $n$ is the data-driven parameter and is the same as the $n$ in (LABEL:eqn_opt), $\Delta n$ is caused by the uncertainty of soft materials, $\theta$ is the bending angle as shown in Fig. 2 (e), $F(p)$ is the equivalent force generated by the input pressure $p$ from the syringe pump, $M_{eq}$ is the equivalent mass of the soft actuator, $C_{n}$ is the damping constant of the soft actuator. The $M_{eq}$ and $C_{n}$ currently are estimated by performing system identification applied to experimentally obtained responses in MATLAB®, which has an average accuracy of approximately $95.3\%$. $K_{n}$ is the spring constant expressed as
@@
where $L_{0}$ is the initial length of the soft actuator, $n$ is the same as the parameter in (LABEL:eqn_opt) and it determined by either experiments or data-driven approach [31], $I_{n}$ is the modified moment of inertia for a large deflection component [29] and it is expressed as
@@
If $n=1$ and $\Delta n=0$, the dynamical equations (9), (10), and (11) become linear equations. Note that the linear model can capture the behavior of the real system within $\theta=0$ to $94\pi/$ 9 $rad$ [31]. In this study, the actuator is aimed to remain within this specific range. Therefore, we set $n=1$ in (9), which yields:
@@
The state-space form is, therefore, written as
@@
where $F(p)$ is represented as $c\cdot p$, $c$ is a constant affected by $a$, $b$, $w$, and $t$ [27]. If we stack the system model of $n_{y}$ fingers, the state-space form becomes
$A_{stk}=\begin{bmatrix}A_{1}&0&0\\ 0&\ddots&0\\ 0&0&A_{n_{y}}\end{bmatrix},B_{stk}=\begin{bmatrix}B_{1}\\ \vdots\\ B_{n_{y}}\end{bmatrix},C_{stk}=\begin{bmatrix}{C_{1}}^{T}\\ \vdots\\ {C_{n_{y}}}^{T}\end{bmatrix}$ (14)
According to (6), we get the system equation in the Laplace domain. Thus, the system model matrix $P(s)^{n_{y}\times 1}$ is described as
${P}(s)=\begin{bmatrix}\frac{c\cdot p/M_{eq}}{s^{2}+(C_{n\_1}/M_{eq})s+K_{n\_1}/M_{eq}}\\ \vdots\\ \frac{c\cdot p/M_{eq}}{s^{2}+(C_{n\_n_{y}}/M_{eq})s+K_{n\_n_{y}}/M_{eq}}\end{bmatrix}$ (15)
where the equivalent mass of the soft actuators is almost the same and the same symbol $M_{eq}$ is used. Although the two fingers share the same dimensional parameters such as height, weight, etc., their $C_{n}$ and $K_{n}$ in (12) are slightly different due to the uncertainty of soft materials.
@@
This nature leads to different motions when a feedback controller is applied. The asynchronized motions further lead to grasping failure[7]. One of the aims of this study is to address such grasping failures by utilizing the algebraic stable inversion approach for SIMO setting, and the experimental results will be presented in Sec. IV-E.
@@ III-B2 Modeling Syringe Pump
The configuration of the syringe pump is visualized in Fig. 2 (f). The dynamical modeling of the syringe pump starts at the linear motor. When the linear motor works, it will move the slider
@@
$\begin{split}Q_{i}=A_{s}v_{s}=\frac{l}{2\pi}\omega_{m}\end{split}$ (18)
@@
where $A_{s}$ is the inner cross-sectional area of the syringe, and $Q_{i}$ is the output air flow rate of the syringe. When $Q_{i}$ is divided by the capacity of the soft actuator, it becomes the pressure-changing rate inside the soft actuator
$\begin{split}\dot{p}=\frac{Q_{i}}{C_{i}}=\frac{A_{s}l}{2\pi C_{i}}\omega_{m}\end{split}$ (19)
@@
The dynamics of the syringe pump is the first-order system. The maximum angular velocity of the motor is 5 $rev/s$. The $C_{i}$ will expand as it is pressurized; however, its effect can be ignored as the input pressure is below 0.1 $MPa$ and the bending angle of the soft actuator is below $32\pi/$ 3 $rad$. The $C_{i}$ here is considered as a constant.
The full model of a single soft actuator is the cascade of the (12) and (19)
$A_{1}=\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&-\frac{K_{n}}{M_{eq}}&-\frac{C_{n}}{M_{eq}}\end{bmatrix},B_{1}=\begin{bmatrix}0\\ 0\\ 1\end{bmatrix},C_{1}=\begin{bmatrix}\frac{c\cdot pA_{s}l}{2\pi C_{i}M_{eq}}\\ 0\\ 0\end{bmatrix}^{T}$ (20)
@@
The resulting systems from $\omega_{m}$ to $\theta$ are third-order and have a pole at the imaginary axis. Hence, the full system model matrix $P(s)^{n_{y}\times 1}$ in Laplace domain can be obtained by referencing (6) and (14)
${P(s)}=\begin{bmatrix}\frac{c\cdot pA_{s}l/2\pi C_{i}M_{eq}}{s^{3}+(C_{n\_1}/M_{eq})s^{2}+(K_{n\_1}/M_{eq})s}\\ \vdots\\ \frac{c\cdot pA_{s}l/2\pi C_{i}M_{eq}}{s^{3}+(C_{n\_n_{y}}/M_{eq})s^{2}+(K_{n\_n_{y}}/M_{eq})s}\end{bmatrix}$ (21)
@@
The full system matrix is causal, and the minimum phase and the $P(s)$ is full rank and invertible.
Even though all finger angles ($\theta$) remain within $[0,4\pi/9]$, equation (20) cannot accurately capture the exact behavior of the soft gripper due to its inherent uncertainty structure [17]. To achieve a more accurate representation, we can modify the model in (12) using a multiplicative uncertainty approach as in [32], defining a set of all possible plants for each finger as follows:
$\Pi:=\{(I+{\Delta}{W_{T}})P~{}|~{}\forall\parallel\Delta\parallel_{\infty}\leq\gamma\}$ (22)
Here, the transfer function $W_{T}\in\mathcal{RH}_{\infty}$ represents the spatial and frequency characteristics of the uncertainty. $\Delta$ denotes any unstructured and unknown yet stable function [32]. A general approach to defining the robustness weight function $W_{T}$ is described below [33]:
$\left|\frac{M_{ik}e^{j\phi_{ik}}}{M_{i}e^{j\phi_{i}}}-1\right|\leq\left|W_{T}(j\omega_{i})\right|,i=1,\ldots,m;k=1,\ldots,n_{r}$ (23)
The magnitude and phase values are assessed over a range of frequencies, denoted as $\omega_{i}$ (ranging from $i=1$ to $m$), and the experiment is repeated $n_{r}$ times. The notation $(M_{ik},\phi_{ik})$ refers to the magnitude-phase measurements corresponding to frequency $\omega_{i}$ and the $k$th experiment iteration, where $k=1$ to $n_{r}$. Similarly, $(M_{i},\phi_{i})$ represents the magnitude-phase pairs for the nominal plant $P(s)$.
The following remark describes an important dependency of the uncertainty in SPAs:
Remark 1.
The single input (underactuated control) may lead to different motions of soft fingers with the same dimensional parameters due to the uncertainty of the soft materials. The deformation curves of some soft materials (Smooth-on Ecoflex series) exhibit high uncertainty when they have a slow deformation rate [16, 34]. In contrast, the curves demonstrate much less uncertainty when their deformation rate is high. Similarly, when a higher pressure changing rate is applied to soft fingers, they show a narrower uncertainty band, align more closely with nominal behaviors, and tend to have consistent motions. This property influences the feasibility and performance of underactuated control of the multi-finger soft gripper (24).
Using the insight from the above remark, equation (22) can be redefined with respect to the operational speed $\omega_{m}$. For simplicity, we consider only two fingers:
$\Pi_{H}(\omega_{m}^{H}):=\Biggl{\{}\Biggl{(}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}+\Delta\overbrace{\begin{bmatrix}W_{T1}&0\\ 0&W_{T2}\end{bmatrix}}^{W_{T}^{H}}\Biggl{)}\begin{bmatrix}P_{1}(s)\\ P_{2}(s)\end{bmatrix}\mid\|\Delta\|_{\infty}\leq\gamma\Biggl{\}}$
$\Pi_{L}(\omega_{m}^{L}):=\Biggl{\{}\Biggl{(}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}+\Delta\underbrace{\begin{bmatrix}W_{T1}&0\\ 0&W_{T2}\end{bmatrix}}_{W_{T}^{L}}\Biggl{)}\begin{bmatrix}P_{1}(s)\\ P_{2}(s)\end{bmatrix}\mid\|\Delta\|_{\infty}\leq\gamma\Biggl{\}}$
such that
$\omega_{m}^{H}>\omega_{m}^{L}\implies\bar{\sigma}(W_{T}^{H}(j\omega_{i}))<\bar{\sigma}(W_{T}^{L}(j\omega_{i})).$
where $\Pi_{H}$ denotes the uncertain plant family with respect to high speed of actuation ($\omega_{m}^{H}$) and $\Pi_{L}$ defines the uncertain plants for low speed of actuation ($\omega_{m}^{L}$). Thus, as the operational speed of the soft actuator increases, the variance in its behavior is reduced.
Figure 3: The block diagram of the proposed controller including the feedforward control and feedback loop.
@@ III-C Controller Design and Analysis
If all the elements in $P(s)$ of (6) are the same (i.e., $P_{1}=P_{2}=...=P_{n_{y}}$), the multiple systems will be coordinated automatically. However, the full system matrix of multiple soft fingers $P(s)$, which has different elements (i.e., $P_{1}\neq P_{2}\neq...\neq P_{n_{y}}$), will be utilized to design the underactuated controller. We can re-formulate (6) for the desired output as
$\begin{bmatrix}Y_{d1}(s)\\ \vdots\\ Y_{d{n_{y}}}(s)\end{bmatrix}=\begin{bmatrix}P_{1}(s)\\ \vdots\\ P_{n_{y}}(s)\end{bmatrix}{U(s)}$ (24)
where $Y_{d}(s)\in\textnormal{Im}_{\mathbb{R}(s)}(P)$. In (24), $Y_{di}(s)$ represents the desired output of $P_{i}(s)$, and $i=1,...,n_{y}$. The stable inversion is composed of the feedforward and feedback loop [21]. The feedforward controller is obtained by solving (24) to get $U(s)$ and the additional feedback loop aims to address the system perturbation of the mechatronic system as shown in Fig. 3.
Some results are introduced and will be implemented to design the controller, including the feedforward control and feedback loop.
@@
(section III-B,  [21]) Let $P(s)$ be non-square ($\textnormal{rank}_{\mathbb{R}(s)}(P)=n_{u}<n_{y}$), then there exists an $P\textsuperscript{\textdagger}(s):=(P^{T}(s)P(s))^{-1}P^{T}(s)$ satisfying $P\textsuperscript{\textdagger}(s)P(s)=I$. Besides, it is defined that $y_{d}(t)$ is the desired system response in the time domain and $y_{d}^{a}(t)$ is the system response by applying an approximate solution $U^{a}(s)$. Thus, an approximate solution $U^{a}(s)$ is defined as
@@
4) $U^{a}(s)\in\mathcal{RH}_{\infty}$
@@
(section III-B, [35]) Let $\omega_{cl}$ denote the bandwidth (BW) of the system. So the $H(s)$ is defined as
$H(j\omega)\approx I\Leftrightarrow~{}\omega\ll\omega_{cl}$
$H(j\omega)\approx 0\Leftrightarrow~{}\omega\gg\omega_{cl}$
$H(j\omega)\not\approx\{0,I\}\Leftrightarrow~{}\omega$ close to $\omega_{cl}$
Remark 2.
The P(s) in (25) contains non-causal elements. To make it applicable to real systems, the $H(s)$ could be loop-shaping synthesis [32] or a low-pass filter [36] to re-shape the P(s). Considering a general case, a system would be minimal or non-minimal. Loop-shaping synthesis is applied to shape the (25). However, if the system is minimal as stated in the Assumption 1 and has a lower order, the low-pass filter, which is relatively more applicable, can be utilized to shape the (25). The order of the low-pass filter depends on the relative order ($l$) of P(s). The low-pass filter takes the form of $\frac{a_{0}}{s^{l}+a_{l-1}s^{l-1}+...+a_{0}}$. The parameters $a_{0},\ldots a_{l-1}$ in the equation are selected to define the cut-off frequency of the low-pass filter.
Remark 3.
An alternative reason to apply $H(s)$ is that using the P(s) in (25) directly in the feedforward or feedback controllers can cause undesired high-frequency excitation. Thus, an appropriate selection of the either cut-off frequency of the low-pass filter or the bandwidth of the resulting complementary sensitivity function of the loop shaping can prevent this excitation. As demonstrated in the experimental part, there is a definite advantage of letting the bandwidth of $H(s)$ be as large as possible so that the better synchronization of multiple fingers be achieved due to the nonlinear nature of the material of soft fingers.
The tracking error (with perturbed term) can be compensated by using the output feedback as displayed in Fig. 3. The $u_{ff}$ is calculated based on the Theorem 2. It is assumed that the output of the real (uncertain) system can be measured such as sensors or observers. Since the nominal system output can be computed with the combined input $u_{c}$, we have the output difference $y_{\Delta}(t)$. With this output difference, we can compensate for the error by the following theorem.